# LLM-Driven Multi-step Translation from C to Rust using Static Analysis

### v2

@@ F.1 Unidiomatic Translation
F.2 Unidiomatic Translation with Feedback
@@ F.4 Idiomatic Verification
This is the idiomatic translation of Rust code from C, the function signature is
@@
{FUNCTION_NAME}_idiomatic
This is the unidiomatic translation of Rust code from C, the function signature is
@@
{FUNCTION_DEPENDENCIES}
Generate the harness for the function atoi_idiomatic with the following code pattern so that it can be tested:
Finish all the TODOs.
You should **NOT** add any dummy implementation of the function or structs, as it will be provided by the verifier:
@@
// TODO: add necessary uses here
// Dont add the definitions of any other functions and structs, they will be provided by the system
fn {FUNCTION_NAME} ({UNIDIOMATIC_ARGS}) -> {UNIDIOMATIC_RETURN} {
// TODO: Add code here to Convert the input to the idiomatic format
let result = fn {FUNCTION_NAME}_idiomatic({IDIOMATIC_ARGS}) -> {IDIOMATIC_RETURN}; // Call the idiomatic function
// TODO: Add code here to Convert the result back to the original format
remove all the TODOs and replace them with the necessary code.
@@
F.5 Failure Reason Analysis

### v3

@@ Appendix I SACTOR Cost Analysis
N.2 Unidiomatic Translation with Feedback
@@
Also output a minimal JSON spec that maps the unidiomatic Rust layout to the idiomatic Rust for the function arguments and return value.
Full JSON Schema for the SPEC (do not output the schema; output only an instance that conforms to it):
‘‘‘json
{_schema_text}
----SPEC----
‘‘‘json
"function_name": "{function.name}",
"fields": [
"u_field": {{
"name": "...",
"type": "...",
"shape": "scalar" | {{"ptr": {{"kind": "slice|cstring|ref", "len_from": "?", "len_const": 1}}}}
"i_field": {{
"name": "...",
"type": "..."
----END SPEC----
Few-shot examples (each with unidiomatic Rust signature, idiomatic Rust signature, and the SPEC):
Example F1 (slice arg):
Unidiomatic Rust:
‘‘‘rust
pub unsafe extern "C" fn sum(xs: *const i32, n: usize) -> i32;
Idiomatic Rust:
‘‘‘rust
pub fn sum(xs: &[i32]) -> i32;
----SPEC----
‘‘‘json
"function_name": "sum",
"fields": [
{{ "u_field": {{"name": "xs", "type": "*const i32", "shape": {{ "ptr": {{ "kind": "slice", "len_from": "n" }} }} }},
"i_field": {{"name": "xs", "type": "&[i32]" }} }},
{{ "u_field": {{"name": "n", "type": "usize", "shape": "scalar" }},
"i_field": {{"name": "xs.len", "type": "usize" }} }}
----END SPEC----
Example F2 (ref out):
Unidiomatic Rust:
‘‘‘rust
pub unsafe extern "C" fn get_value(out_value: *mut i32);
Idiomatic Rust:
‘‘‘rust
pub fn get_value() -> i32;
----SPEC----
‘‘‘json
"function_name": "get_value",
"fields": [
{{ "u_field": {{"name": "out_value", "type": "*mut i32", "shape": {{ "ptr": {{ "kind": "ref" }} }} }},
"i_field": {{"name": "ret", "type": "i32" }} }}
----END SPEC----
Example F3 (nullable cstring maps to Option):
Unidiomatic Rust:
‘‘‘rust
pub unsafe extern "C" fn set_name(name: *const libc::c_char);
Idiomatic Rust:
‘‘‘rust
pub fn set_name(name: Option<&str>);
----SPEC----
‘‘‘json
"function_name": "set_name",
"fields": [
{{ "u_field": {{"name": "name", "type": "*const c_char", "shape": {{ "ptr": {{ "kind": "cstring", "null": "nullable" }} }} }},
"i_field": {{"name": "name", "type": "Option<&str>" }} }}
----END SPEC----
@@
We have an initial spec-driven harness with TODOs. Finish all TODOs and ensure it compiles.
Idiomatic signature:
@@
pub fn compute_idiomatic(
x: i32,
name: &str,
data: &[u8],
meta: HashMap<String, String>,
) -> i32;;
Unidiomatic signature:
@@
pub unsafe extern "C" fn compute(x: i32, name: *const libc::c_char, data: *const u8, len: usize, meta: *const libc::c_char) -> i32;;
Current harness:
@@
pub unsafe extern "C" fn compute(x: i32, name: *const libc::c_char, data: *const u8, len: usize, meta: *const libc::c_char) -> i32
// Arg name’: borrowed C string at name
let name_str = if !name.is_null() {
unsafe { std::ffi::CStr::from_ptr(name) }.to_string_lossy().into_owned()
} else {
String::new()
// Arg data’: slice from data with len len as usize
let data_len = len as usize;
let data_len_non_null = if data.is_null() { 0 } else { data_len };
let data: &[u8] = if data_len_non_null == 0 {
} else {
unsafe { std::slice::from_raw_parts(data as *const u8, data_len_non_null) }
// TODO: param meta of type HashMap < String , String >: unsupported mapping
let __ret = compute_idiomatic(x, &name_str, data, /* TODO param meta */);
return __ret;
Output only the final function in this format:
@@
N.5 Failure Reason Analysis